HETAV PATEL

Grade 11,

if the P.E of hydrogen electron is -3.02 eV then in which excited level is the electron present? a.) 1 b.) 2 c.) 3 d.) 4

if the P.E of hydrogen electron is -3.02 eV then in which excited level is the electron present?
a.) 1   b.) 2   c.) 3   d.) 4 

Grade:11

1 Answers

Sourav Maji
6 years ago
P.E=1/2 Total Energy Then, Total energy =1.51evNow, use the formulaE(n) = -13.6*(Z/n)^2For hydrogen Z=1Now, -1.51=-13.6/n^2 =>n^2=9(approx) n=3So the electron will be present in 3rd orbit or 2nd excited state. So, answer b) is correct.